# Residually finite groups with uniformly almost flat quotients

### v1

@@ 1. Introduction
Theorem 1.3 (Khukhro–Valette [16, Theorem 3]).
@@
For smaller values of $\alpha$ the equivalence between being strongly residually $\alpha$-almost flat, having $\alpha$-almost flat quotients, and having $\alpha$-almost flat coset spaces breaks down. This is most striking for strong residual almost flatness, where Khukhro and Valette show that for any $\alpha<1$, every finitely generated residually finite group that surjects onto $\mathbb{Z}$ is strongly residually $\alpha$-almost flat [16, Proposition 5]. In our first result, we show that moreover having $\alpha$-almost flat quotients is not equivalent to having $\alpha$-almost flat coset spaces, as follows.
@@
Our approach to Theorem 1.5 is inspired by the proof of Wolf’s theorem that a polycyclic group has polynomial growth if and only if it virtually nilpotent as presented for example in [4, Chapter 14], but is considerably more complicated. One difficulty lies in actually constructing, in a polycyclic group of superpolynomial growth, an infinite sequence of finite-index subgroups whose coset spaces violate the uniformly $\alpha$-almost flat condition. This turns out to require various number-theoretic facts, as well as a result of Beruillard and the second author [2, Theorem 4.1] that in turn relies on Breuillard, Green and Tao’s deep refinement of Gromov’s theorem. Another issue is that the proof of Wolf’s theorem uses the fact that having polynomial growth passes to finitely generated subgroups, but a priori we do not have the analogous statement that having uniformly $\alpha$-almost flat coset spaces passes to finitely generated subgroups.
@@
In his forthcoming PhD thesis, the first author will show that the coset-space version of 1.7 holds for all finitely generated residually finite $-\mathbb{Z}$ groups. In the special cases of the Baumslag–Solitar groups $\mathrm{BS}(1,k)$ with $k\geqslant 2$ and wreath products $A\wr\mathbb{Z}$ with $A$ abelian, he will show that the full conjecture holds. Since wreath products $H\wr\mathbb{Z}$ with $H$ non-abelian are not residually finite [10, Theorem 3.2], this completely settles 1.7 for groups of this form.
@@ Acknowledgements
We are indebted to James Rawson, who investigated the property of having uniformly almost flat quotients under the supervision of the second author through the Summer Research in Maths programme at the University of Cambridge. In particular, we thank him for the suggestion to consider groups with uniformly almost flat coset spaces in addition to those with uniformly almost flat quotients. The first author thanks Yves Cornulier, and the second author thanks Ana Khukhro and Alain Valette, for helpful conversations. We also thank Alain Valette for comments on a draft.
@@ 2. Basic properties of uniform almost flatness
By the definition of diameter, $S^{\operatorname{diam}_{S}(G/H)}$ contains a set of coset representations of $H$ in $G$, call this set $C$. It is shown in the proof of [11, Lemma 7.2.2] that $S^{m}\cap H\subseteq(CSC^{-1}\cap H)^{m}$ for all $m\in\mathbb{N}$, and hence in particular that $CSC^{-1}\cap H\subseteq S^{2\operatorname{diam}_{S}(G/H)+1}$ generates $H$ with diameter at most $\operatorname{diam}_{S}(G)$. ∎
@@
For the first inequality, let $h\in H$, then $h\in s_{1}\ldots s_{m}K$ for some $m\leqslant\operatorname{diam}_{S}(G/K)$ and $s_{i}\in S$. Since $K\subseteq H$, $s_{1}\ldots s_{m}\in H$. Let $T$ be the generating set for $H$ defined in Lemma 2.3. Applying [11, Lemma 7.2.2] once more as in the proof of the previous lemma, we can write $s_{1}\ldots s_{m}$ as $t_{1}\ldots t_{m}$ for some $t_{i}\in T$. Hence $h\in t_{1}\ldots t_{m}K$. For the second inequality, by Lemma 2.3, $S^{3\operatorname{diam}_{S}(G/K)}\cap H$ generates $H$. Therefore, ($S^{3\operatorname{diam}_{S}(G/H)}\cap H)^{\operatorname{diam}_{S}(G/K)}$ contains a set of coset representative for $K$ in $H$. Hence $H\subseteq S^{3\operatorname{diam}_{S}(G/H)\operatorname{diam}_{S}(G/K)}K$. Hence, $G=S^{\operatorname{diam}(G/H)}H=S^{\operatorname{diam}(G/H)}(S^{3\operatorname{diam}_{S}(G/H)\operatorname{diam}_{S}(G/K)}K)=S^{4\operatorname{diam}_{S}(G/H)\operatorname{diam}_{S}(G/K)}K$. ∎
@@ 4. Khukhro and Valette’s results
The implication (2$\implies$(1) is trivial. The implication (3$\implies$(2) is proved in [16, Theorem 4]; indeed, it is shown there that for any finite generating set $S$ there exists a decreasing sequence $(N_{k})_{k\in\mathbb{N}}$ of finite-index normal subgroups with trivial intersection and $\varepsilon,\alpha\in(0,1)$ such that $\operatorname{diam}_{S}(G/N_{k})\geqslant\varepsilon[G:N_{k}]^{\alpha}$ for all $k$. It remains to prove (1$\implies$(3), which we do following the proof of [2, Corollary 1.7].
@@ 5. Residually nilpotent groups
Following [20], for each $k\in\mathbb{N}$ we also define the generalised commutator subgroup $\overline{G}_{k}$ by $\overline{G}_{k}=\{g\in G:(\exists n\in\mathbb{N})(g^{n}\in G_{k})\}$. The generalised commutator subgroups are all characteristic in $G$, each $\overline{G}_{i}/\overline{G}_{i+1}$ is torsion-free abelian, and if $G$ is finitely generated then each $\overline{G}_{k}$ contains $G_{k}$ as a finite-index subgroup [20, Lemma 4.1]. Moreover, we have $[\overline{G}_{i},\overline{G}_{j}]\leqslant\overline{G}_{i+j}$ for all $i,j\in\mathbb{N}$ [20, Lemma 4.5].
If $N$ is a simply connected nilpotent Lie group with Lie algebra $\mathfrak{n}$ then there exist mutually inverse diffeomorphisms $\exp:\mathfrak{n}\to N$ and $\log:N\to\mathfrak{n}$ [3, Theorem 1.2.1]. The lower central series $\mathfrak{n}=\mathfrak{n}_{1}\geqslant\mathfrak{n}_{2}\geqslant\cdots$ of $\mathfrak{n}$ is defined recursively by $\mathfrak{n}_{i+1}=[\mathfrak{n},\mathfrak{n}_{i}]=\mathrm{span}_{\mathbb{R}}\{[x,y]:x\in\mathfrak{n},y\in\mathfrak{n}_{i}\}$. The terms $N_{i}$ of the lower central series of $N$ are Lie subgroups and $\exp N_{i}=\mathfrak{n}_{i}$ for each $i$ [12, Theorem XII.3.1]. Simply connected nilpotent Lie groups are uniquely divisible: for every $x\in N$ and $k\in\mathbb{N}$ there exists a unique $y\in N$ such that $y^{k}=x$.
@@
This is standard and implicit in [3], so we provide only brief details of how it may be read out of that source. The normal subgroups $N_{i}$ are rational by [3, Theorem 5.1.8 & Corollary 5.2.2], so the existence of a basis satisfying conditions (i)–(iii) follows from [3, Corollary 5.3.2 & Proposition 1.2.7]. The proof of [3, Theorem 5.1.8 (a)] then shows (as stated in the second sentence of the proof) that the structure constants are rational. ∎
@@
Let $d$ be the lowest common denominator of the rationals $c_{ijk}$. For each $i$ set $f_{i}=de_{i}$, noting that $[f_{i},f_{j}]=\sum_{k=a+1}^{h}dc_{ijk}f_{k}$ for all $i,j$. Let us emphasise that the coefficients $dc_{ijk}$ in this sum are integers. Given $m\in\mathbb{N}$, we therefore have $[m^{b(i)}f_{i},m^{b(j)}f_{j}]=\sum_{k=a+1}^{h}dc_{ijk}m^{b(i)+b(j)}f_{k}$ for all $i,j$. Writing $\Lambda$ for the additive subgroup of $\mathfrak{n}$ spanned by $m^{b(i)}f_{i}$, it follows that $[\Lambda,\Lambda]\subseteq\Lambda$, so [27, Lemma 4.3] implies that there exists $r$ depending only on $c$ such that $\exp(r\Lambda)$ is a subgroup of $N$, and [26, Corollary 3.11] then implies that $\exp(r\Lambda)=\{x_{1}^{drm^{b(1)}\ell_{1}}\cdots x_{h}^{drm^{b(h)}\ell_{h}}:\ell_{i}\in\mathbb{Z}\}$, giving (iii). ∎
@@ 6.2. Some standard results from number theory
It is well-known that every non-constant polynomial in $\mathbb{Z}[X]$ splits completely over infinitely many fields of prime order (see e.g. [6, Corollary on p. 258]). For every such prime greater than $P(0)$, it is clear that $P(x)$ has no zero roots over $\mathbb{F}_{p}$. ∎
@@
Let $R$ be an integral domain, we will also make use of the function $res:R[x]\times R[x]\rightarrow R$, the resultant of two polynomials. The resultant $res(f,g)$ of $f$ and $g$ is defined as the determinant of their Sylvester matrix, which is a matrix whose entries are either zero or the coefficients of $f$ and $g$. One consequence of the definition is that, given $f$ , $g\in\mathbb{Z}[x]$ and a prime $p$, we have $\operatorname{res}(f\bmod{p},g\bmod{p})=\operatorname{res}(f,g)\bmod{p}$. We refer the readers to [5, §12] for the relevant details. Here is an important property of this function we will use later.
Theorem 6.7 ([5, §12]).
@@ 6.3. The multiplicative order of the eigenvalues
Given a commutative ring $R$, every monic polynomial $P(X)\in R[X]$ is a characteristic polynomial of some matrix with entries in $R$; indeed, one such matrix is thc companion matrix, constructed in [13, §3.3].
@@ 6.4. $\mathbb{Z}^{n}$ by $\mathbb{Z}$ groups
We will be using the same set-up as [4, §14]. The semidirect product $\mathbb{Z}^{n}\rtimes_{\Phi}\mathbb{Z}$ is defined by a homomorphism $\Phi:\mathbb{Z}\rightarrow Aut(\mathbb{Z}^{n})=\operatorname{GL}(n,\mathbb{Z})$, and $\phi$ is entirely determined by the automorphism $\theta\coloneqq\phi(1)$, which can be represented by a matrix $M\in\operatorname{GL}(n,\mathbb{Z})$. We can therefore denote the semidirect product $\mathbb{Z}^{n}\rtimes_{\phi}\mathbb{Z}$ as $\mathbb{Z}^{n}\rtimes_{M}\mathbb{Z}$.
@@
The equivalence (i) $\Leftrightarrow$ (ii) is proved in [4, Proposition 14.1]. The implication (ii) $\Rightarrow$ (iii) is proved in Lemma 2.2. The implication (iii) $\Rightarrow$ (i) is proved in Proposition 6.17. The implication (i) $\Rightarrow$ (iv) is proven in [4, Lemma 13.28]. ∎
@@ 7. $-\mathbb{Z}$ groups
Recall that the set of all finite order elements forms a characteristic subgroup of $G$, we will denote it by $\operatorname{Tor}G$. Suppose $G$ is $k$-step nilpotent, i.e. $G$ has the lower central series $G=G_{1}\geqslant G_{2}\geqslant\ldots G_{k+1}=\{\mathrm{id}_{G}\},$ where $G_{i+1}=[G_{i},G]\operatorname{char}G$. Let $\phi_{i}\coloneqq\phi|_{G_{i}}$ be the restriction of $\phi$ on $G_{i}$. Then $\phi_{i}$ projects to an automorphism $\theta_{i}$ on the finitely generated abelian group $B_{i}\coloneqq\frac{G_{i}}{G_{i+1}}$. The automorphism $\theta_{i}$ then projects to an automorphism $\bar{\theta}_{i}$ on the torsion free quotient $\frac{B_{i}}{\operatorname{Tor}{B_{i}}}\cong\mathbb{Z}^{m_{i}}$. Fix a basis for each $\frac{B_{i}}{\operatorname{Tor}{B_{i}}}\cong\mathbb{Z}^{m_{i}}$, then each automorphism $\bar{\theta}_{i}$ is corresponds to a matrix $M_{i}\in\operatorname{GL}(m_{i},\mathbb{Z})$.
In the proof of characterising nilpotent by $\mathbb{Z}$ groups in terms of their growth rate [4, Proposition 14.28], there are two cases to consider, whether or not all of $M_{i}$ have eigenvalues of absolute value $1$. It turns out that it is enough to examine the first matrix $M_{1}$.
@@
The implication(i) $\Rightarrow$ (ii) is proved in Lemma 7.1. The implication (ii) $\Rightarrow$ (i) is clear. The equivalence (ii) $\Leftrightarrow$ (iii) is proved in [4, Proposition 14.28]. The implication (iii) $\Rightarrow$ (iv) is proved in Lemma 2.2. The implication (iv) $\Rightarrow$ (i) is proved in Proposition 7.5. Finally, the implication (i) $\Leftrightarrow$ (v) is proved in Proposition 6.17. ∎

### v2

@@ 1. Introduction
Theorem 1.3 (Khukhro–Valette [22, Theorem 3]).
@@
For smaller values of $\alpha$ the equivalence between being strongly residually $\alpha$-almost flat, having $\alpha$-almost flat quotients, and having $\alpha$-almost flat coset spaces breaks down. This is most striking for strong residual almost flatness, where Khukhro and Valette show that for any $\alpha<1$, every finitely generated residually finite group that surjects onto $\mathbb{Z}$ is strongly residually $\alpha$-almost flat [22, Proposition 5]. In our first result, we show that moreover having $\alpha$-almost flat quotients is not equivalent to having $\alpha$-almost flat coset spaces, as follows.
@@
The finitely generated residually torsion-free nilpotent groups include many well-known examples of groups, such as free groups [3, Theorem 1], surface groups [3, Theorem 5], right-angled Artin groups [9], the Torelli group of a right-angled Artin group [34], pure braid groups [1], the Torelli subgroup of the mapping class group [2] and many more (e.g. see [4, 5]).
Our approach to Theorem 1.5 is inspired by the proof of Wolf’s theorem that a polycyclic group has polynomial growth if and only if it virtually nilpotent as presented for example in [10, Chapter 14], but is considerably more complicated. One difficulty lies in actually constructing, in a polycyclic group of superpolynomial growth, an infinite sequence of finite-index subgroups whose coset spaces violate the uniformly $\alpha$-almost flat condition. This turns out to require various number-theoretic facts, as well as a result of Beruillard and the second author [7, Theorem 4.1] that in turn relies on Breuillard, Green and Tao’s deep refinement of Gromov’s theorem. Another issue is that the proof of Wolf’s theorem uses the fact that having polynomial growth passes to finitely generated subgroups, but a priori we do not have the analogous statement that having uniformly $\alpha$-almost flat coset spaces passes to finitely generated subgroups.
@@
In his forthcoming PhD thesis, the first author will show that the coset-space version of 1.7 holds for all finitely generated residually finite $-\mathbb{Z}$ groups. In the special cases of the Baumslag–Solitar groups $\mathrm{BS}(1,k)$ with $k\geqslant 2$ and wreath products $A\wr\mathbb{Z}$ with $A$ abelian, he will show that the full conjecture holds. Since wreath products $H\wr\mathbb{Z}$ with $H$ non-abelian are not residually finite [16, Theorem 3.2], this completely settles 1.7 for groups of this form.
@@ Acknowledgements
We are indebted to James Rawson, who investigated the property of having uniformly almost flat quotients under the supervision of the second author through the Summer Research in Maths programme at the University of Cambridge. In particular, we thank him for the suggestion to consider groups with uniformly almost flat coset spaces in addition to those with uniformly almost flat quotients. The first author thanks Yves Cornulier, and the second author thanks Ana Khukhro and Alain Valette, for helpful conversations. We also thank Alain Valette for comments on a draft, and Mark Hagen for help with the references.
@@ 2. Basic properties of uniform almost flatness
By the definition of diameter, $S^{\operatorname{diam}_{S}(G/H)}$ contains a set of coset representations of $H$ in $G$, call this set $C$. It is shown in the proof of [17, Lemma 7.2.2] that $S^{m}\cap H\subseteq(CSC^{-1}\cap H)^{m}$ for all $m\in\mathbb{N}$, and hence in particular that $CSC^{-1}\cap H\subseteq S^{2\operatorname{diam}_{S}(G/H)+1}$ generates $H$ with diameter at most $\operatorname{diam}_{S}(G)$. ∎
@@
For the first inequality, let $h\in H$, then $h\in s_{1}\ldots s_{m}K$ for some $m\leqslant\operatorname{diam}_{S}(G/K)$ and $s_{i}\in S$. Since $K\subseteq H$, $s_{1}\ldots s_{m}\in H$. Let $T$ be the generating set for $H$ defined in Lemma 2.3. Applying [17, Lemma 7.2.2] once more as in the proof of the previous lemma, we can write $s_{1}\ldots s_{m}$ as $t_{1}\ldots t_{m}$ for some $t_{i}\in T$. Hence $h\in t_{1}\ldots t_{m}K$. For the second inequality, by Lemma 2.3, $S^{3\operatorname{diam}_{S}(G/K)}\cap H$ generates $H$. Therefore, ($S^{3\operatorname{diam}_{S}(G/H)}\cap H)^{\operatorname{diam}_{S}(G/K)}$ contains a set of coset representative for $K$ in $H$. Hence $H\subseteq S^{3\operatorname{diam}_{S}(G/H)\operatorname{diam}_{S}(G/K)}K$. Hence, $G=S^{\operatorname{diam}(G/H)}H=S^{\operatorname{diam}(G/H)}(S^{3\operatorname{diam}_{S}(G/H)\operatorname{diam}_{S}(G/K)}K)=S^{4\operatorname{diam}_{S}(G/H)\operatorname{diam}_{S}(G/K)}K$. ∎
@@ 4. Khukhro and Valette’s results
The implication (2$\implies$(1) is trivial. The implication (3$\implies$(2) is proved in [22, Theorem 4]; indeed, it is shown there that for any finite generating set $S$ there exists a decreasing sequence $(N_{k})_{k\in\mathbb{N}}$ of finite-index normal subgroups with trivial intersection and $\varepsilon,\alpha\in(0,1)$ such that $\operatorname{diam}_{S}(G/N_{k})\geqslant\varepsilon[G:N_{k}]^{\alpha}$ for all $k$. It remains to prove (1$\implies$(3), which we do following the proof of [7, Corollary 1.7].
@@ 5. Residually nilpotent groups
Following [26], for each $k\in\mathbb{N}$ we also define the generalised commutator subgroup $\overline{G}_{k}$ by $\overline{G}_{k}=\{g\in G:(\exists n\in\mathbb{N})(g^{n}\in G_{k})\}$. The generalised commutator subgroups are all characteristic in $G$, each $\overline{G}_{i}/\overline{G}_{i+1}$ is torsion-free abelian, and if $G$ is finitely generated then each $\overline{G}_{k}$ contains $G_{k}$ as a finite-index subgroup [26, Lemma 4.1]. Moreover, we have $[\overline{G}_{i},\overline{G}_{j}]\leqslant\overline{G}_{i+j}$ for all $i,j\in\mathbb{N}$ [26, Lemma 4.5].
If $N$ is a simply connected nilpotent Lie group with Lie algebra $\mathfrak{n}$ then there exist mutually inverse diffeomorphisms $\exp:\mathfrak{n}\to N$ and $\log:N\to\mathfrak{n}$ [8, Theorem 1.2.1]. The lower central series $\mathfrak{n}=\mathfrak{n}_{1}\geqslant\mathfrak{n}_{2}\geqslant\cdots$ of $\mathfrak{n}$ is defined recursively by $\mathfrak{n}_{i+1}=[\mathfrak{n},\mathfrak{n}_{i}]=\mathrm{span}_{\mathbb{R}}\{[x,y]:x\in\mathfrak{n},y\in\mathfrak{n}_{i}\}$. The terms $N_{i}$ of the lower central series of $N$ are Lie subgroups and $\exp N_{i}=\mathfrak{n}_{i}$ for each $i$ [18, Theorem XII.3.1]. Simply connected nilpotent Lie groups are uniquely divisible: for every $x\in N$ and $k\in\mathbb{N}$ there exists a unique $y\in N$ such that $y^{k}=x$.
@@
This is standard and implicit in [8], so we provide only brief details of how it may be read out of that source. The normal subgroups $N_{i}$ are rational by [8, Theorem 5.1.8 & Corollary 5.2.2], so the existence of a basis satisfying conditions (i)–(iii) follows from [8, Corollary 5.3.2 & Proposition 1.2.7]. The proof of [8, Theorem 5.1.8 (a)] then shows (as stated in the second sentence of the proof) that the structure constants are rational. ∎
@@
Let $d$ be the lowest common denominator of the rationals $c_{ijk}$. For each $i$ set $f_{i}=de_{i}$, noting that $[f_{i},f_{j}]=\sum_{k=a+1}^{h}dc_{ijk}f_{k}$ for all $i,j$. Let us emphasise that the coefficients $dc_{ijk}$ in this sum are integers. Given $m\in\mathbb{N}$, we therefore have $[m^{b(i)}f_{i},m^{b(j)}f_{j}]=\sum_{k=a+1}^{h}dc_{ijk}m^{b(i)+b(j)}f_{k}$ for all $i,j$. Writing $\Lambda$ for the additive subgroup of $\mathfrak{n}$ spanned by $m^{b(i)}f_{i}$, it follows that $[\Lambda,\Lambda]\subseteq\Lambda$, so [33, Lemma 4.3] implies that there exists $r$ depending only on $c$ such that $\exp(r\Lambda)$ is a subgroup of $N$, and [32, Corollary 3.11] then implies that $\exp(r\Lambda)=\{x_{1}^{drm^{b(1)}\ell_{1}}\cdots x_{h}^{drm^{b(h)}\ell_{h}}:\ell_{i}\in\mathbb{Z}\}$, giving (iii). ∎
@@ 6.2. Some standard results from number theory
It is well-known that every non-constant polynomial in $\mathbb{Z}[X]$ splits completely over infinitely many fields of prime order (see e.g. [12, Corollary on p. 258]). For every such prime greater than $P(0)$, it is clear that $P(x)$ has no zero roots over $\mathbb{F}_{p}$. ∎
@@
Let $R$ be an integral domain, we will also make use of the function $res:R[x]\times R[x]\rightarrow R$, the resultant of two polynomials. The resultant $res(f,g)$ of $f$ and $g$ is defined as the determinant of their Sylvester matrix, which is a matrix whose entries are either zero or the coefficients of $f$ and $g$. One consequence of the definition is that, given $f$ , $g\in\mathbb{Z}[x]$ and a prime $p$, we have $\operatorname{res}(f\bmod{p},g\bmod{p})=\operatorname{res}(f,g)\bmod{p}$. We refer the readers to [11, §12] for the relevant details. Here is an important property of this function we will use later.
Theorem 6.7 ([11, §12]).
@@ 6.3. The multiplicative order of the eigenvalues
Given a commutative ring $R$, every monic polynomial $P(X)\in R[X]$ is a characteristic polynomial of some matrix with entries in $R$; indeed, one such matrix is thc companion matrix, constructed in [19, §3.3].
@@ 6.4. $\mathbb{Z}^{n}$ by $\mathbb{Z}$ groups
We will be using the same set-up as [10, §14]. The semidirect product $\mathbb{Z}^{n}\rtimes_{\Phi}\mathbb{Z}$ is defined by a homomorphism $\Phi:\mathbb{Z}\rightarrow Aut(\mathbb{Z}^{n})=\operatorname{GL}(n,\mathbb{Z})$, and $\phi$ is entirely determined by the automorphism $\theta\coloneqq\phi(1)$, which can be represented by a matrix $M\in\operatorname{GL}(n,\mathbb{Z})$. We can therefore denote the semidirect product $\mathbb{Z}^{n}\rtimes_{\phi}\mathbb{Z}$ as $\mathbb{Z}^{n}\rtimes_{M}\mathbb{Z}$.
@@
The equivalence (i) $\Leftrightarrow$ (ii) is proved in [10, Proposition 14.1]. The implication (ii) $\Rightarrow$ (iii) is proved in Lemma 2.2. The implication (iii) $\Rightarrow$ (i) is proved in Proposition 6.17. The implication (i) $\Rightarrow$ (iv) is proven in [10, Lemma 13.28]. ∎
@@ 7. $-\mathbb{Z}$ groups
Recall that the set of all finite order elements forms a characteristic subgroup of $G$, we will denote it by $\operatorname{Tor}G$. Suppose $G$ is $k$-step nilpotent, i.e. $G$ has the lower central series $G=G_{1}\geqslant G_{2}\geqslant\ldots G_{k+1}=\{\mathrm{id}_{G}\},$ where $G_{i+1}=[G_{i},G]$. Let $\phi_{i}\coloneqq\phi|_{G_{i}}$ be the restriction of $\phi$ on $G_{i}$. Then $\phi_{i}$ projects to an automorphism $\theta_{i}$ on the finitely generated abelian group $B_{i}\coloneqq\frac{G_{i}}{G_{i+1}}$. The automorphism $\theta_{i}$ then projects to an automorphism $\bar{\theta}_{i}$ on the torsion free quotient $\frac{B_{i}}{\operatorname{Tor}{B_{i}}}\cong\mathbb{Z}^{m_{i}}$. Fix a basis for each $\frac{B_{i}}{\operatorname{Tor}{B_{i}}}\cong\mathbb{Z}^{m_{i}}$, then each automorphism $\bar{\theta}_{i}$ is corresponds to a matrix $M_{i}\in\operatorname{GL}(m_{i},\mathbb{Z})$.
In the proof of characterising nilpotent by $\mathbb{Z}$ groups in terms of their growth rate [10, Proposition 14.28], there are two cases to consider, whether or not all of $M_{i}$ have eigenvalues of absolute value $1$. It turns out that it is enough to examine the first matrix $M_{1}$.
@@
The implication(i) $\Rightarrow$ (ii) is proved in Lemma 7.1. The implication (ii) $\Rightarrow$ (i) is clear. The equivalence (ii) $\Leftrightarrow$ (iii) is proved in [10, Proposition 14.28]. The implication (iii) $\Rightarrow$ (iv) is proved in Lemma 2.2. The implication (iv) $\Rightarrow$ (i) is proved in Proposition 7.5. Finally, the implication (i) $\Leftrightarrow$ (v) is proved in Proposition 6.17. ∎